I'm trying to help my niece with her homework and I can't remember at all how to solve this problem.
She's trying to solve this linear equation or system. She says she's supposed to get the answer by using the 3y?
2x-3y=-7 (negative 7)
3x+y=-5 (negative 5)
I have the answer for her but would really like to know how to solve this and to show her the correct way.
Thank you!

Answer by ankor@dixie-net.com(22740)

2x - 3y = -7
3x + y = -5
They probably want you to use elimination, by multiplying the 2nd equation by
3 and adding the two equations, you will eliminate y and easily can find x
:
2x - 3y = -7
9x + 3y = -15 (multiplied by 3)
----------------
11x + 0 = -22
Divide both sides by 11, and you have
x = -2
:
You can now find y by substituting -2 for x in the 2nd equation
3(-2) + y = -5
-6 + y = -5
Add 6 to both sides
y = -5 + 6
y = +1
:
Check your solutions by substitution in the 1st equation
2x - 3y = -7
2(-2) - 3(1) = -7
-4 - 3 = -7; confirms our solutions of x = -2, y = 1